### v1

@@ Abstract
Turbulent open channel flow is known to feature a multi-layer structure near the free surface. In the present work we employ direct numerical simulations considering Reynolds numbers up to $\mathrm{Re}_{\tau}=900$ and domain sizes large enough ($L_{x}=12\pi h$, $L_{z}=4\pi h$) to faithfully capture the effect of very-large-scale motions in order to test the proposed scaling laws and ultimately answer the question: How far does the influence of the free surface extend? In the region near the free surface, where fluctuation intensities of velocity and vorticity become highly anisotropic, we observe the previously documented triple-layer structure, consisting of a wall-normal velocity damping layer that scales with the channel height $h$, and two sublayers that scale with the near-surface viscous length scale $\ell_{V}=\mathrm{Re}_{b}^{-1/2}h$ and with the Kolmogorov length scale $\ell_{K}=\mathrm{Re}_{b}^{-3/4}h$, respectively. The scaling laws previously proposed by Calmet and Magnaudet [J. Fluid. Mech. 474, 355–378 (2003)] are found to hold with the following exceptions. The thin layer, where the intensity of surface-parallel components of the vorticity rapidly decreases to zero, is here found to scale with the Kolmogorov length scale $\ell_{K}$ rather than with the near-surface viscous scale $\ell_{V}$. In addition, we argue that the Kolmogorov length scale is the relevant scale for the mean velocity gradient near the free surface. Both the mean velocity gradient and the fluctuation intensity of the surface-parallel component of vorticity decay to zero in the Kolmogorov sublayer $\delta_{K}\approx 20\ell_{K}$. On the other hand, the layer, where the wall-normal turbulence intensity decreases linearly to zero near the free surface, scales with $\ell_{V}$ rather than $\ell_{K}$ as suggested by Calmet and Magnaudet. The corresponding near-surface viscous sublayer measures $\delta_{V}\approx\ell_{V}$. Importantly, the streamwise turbulence intensity profile for $\mathrm{Re}_{\tau}\geq 400$ suggests that the influence of the free-slip boundary penetrates essentially all the way down to the solid wall through the appearance of enhanced very-large-scale motions ($\delta_{SIL}\approx h$). In contrast, the layer where the surface-normal turbulence intensity is damped to zero is restricted to the free surface ($\delta_{NVD}\approx 0.3h$). As a consequence, the partitioning of the surface-influenced region has to be expanded to a four-layer structure that spans the entire channel height $h$.
@@ 1 Introduction
Plane Poiseuille flow, also known as closed channel flow (CCF), is one of the most studied canonical flows by means of direct numerical simulations [DNSs, 22, 27, 15, 24, 33]. The numerical domain is defined by doubly-periodic boundaries in the stream- and spanwise directions, and impermeable no-slip walls at the bottom and the top. Conversely, less attention has been paid to open channel flow (OCF), where one of the no-slip walls is replaced by a free-slip plane, despite its direct relevance to environmental flows. As noted by Pinelli et al. [37], OCF is of particular interest when studying the mass transfer across gas-liquid interfaces in oceans or engineering applications. One of the defining characteristics of OCF is the transfer of kinetic energy from the vertical to the surface-parallel velocity components due to the damping effect of the free surface [39, 25, 17, 23, 35]. While the inter-component kinetic energy transfer is dominated by the small scales in the vicinity of the free surface, Pinelli et al. [37] showed that low-speed large- and very-large-scale motions ((V)LSMs) are much more correlated with the regions of small-scale vortices than high-speed (V)LSMs. Hence, the spatial distribution of the small scales in the vicinity of the free surface is modulated by the large scales. Most of the early OCF DNSs were restricted to small friction Reynolds numbers $\mathrm{Re}_{\tau}=u_{\tau}h/\nu$, based on friction velocity $u_{\tau}$, channel height $h$ and kinematic viscosity $\nu$, and/or small computational domains. Recently, Yao et al. [41] compared CCF and OCF by means of DNS data in computational domains of $L_{x}\times L_{z}=8\pi h\times 4\pi h$ for Reynolds numbers up to $\mathrm{Re}_{\tau}=2000$. In agreement with experiments [9, 12] and previous DNSs [3], they reported that VLSMs—which are a common feature of wall-bounded turbulent flows—are more energetic and larger in OCF than in CCF. Generally, VLSMs become more energetic with increasing Reynolds number [3, 12, 41], leading to the failure of wall scaling of the streamwise intensity in CCF [15]. As resported by [15] for CCF, the streamwise turbulence intensity scales neither in wall nor in bulk units. To compare wall-attached motions between OCF and CCF, Gong et al. [13] performed DNSs in extremely large domains of length $48\pi h$ for $\mathrm{Re}_{\tau}=550$ and length $24\pi h$ for $\mathrm{Re}_{\tau}=1000$. They found that wall-attached motions contribute more to the turbulent intensity in OCF than in CCF. In Bauer et al. [7], the current authors carried out DNSs at friction Reynolds numbers up to $\mathrm{Re}_{\tau}\approx 900$ in large computational domains up to $L_{x}\times L_{z}=12\pi h\times 4\pi h$ and found that unlike CCF, the streamwise turbulence intensity in OCF scales with the bulk velocity for $\mathrm{Re}_{\tau}\gtrsim 400$. The additional streamwise kinetic energy in OCF was attributed to larger and more intense VLSMs compared to CCF.
@@
CM03 first deduced an integral length scale $L_{\infty}$ using the assumption of high-Reynolds-number equilibrium between kinetic energy transfer and dissipation rate, $\varepsilon\sim u_{iso}^{3}/\ell$, (where $u_{iso}=(2k/3)^{1/2}$ and $\ell$ is the characteristic scale of the energetic eddies) and the relation $L_{\infty}=\ell/2$ which Hunt and Graham [14] found to be valid for homogeneous-isotropic turbulence. Subsequently, CM03 performed a large-eddy simulation (LES) of OCF at $\mathrm{Re}_{\tau}=1280$ and compared their results to these scaling relations. Based on their data, CM03 estimated the thickness of the so-called “surface-influenced layer” as $\delta_{SIL}\approx L_{\infty}\approx 0.2h$, in agreement with the rapid distortion theory of Hunt and Graham [14]. CM03 further categorized the surface-influenced layer into the following sublayers in the near-surface region: a viscous sublayer (of wall-normal extent $\delta_{V}$) which scales with $\mathrm{Re}_{\infty}^{-1/2}L_{\infty}$, and a Kolmogorov sublayer (of extent $\delta_{K}$) which scales with $\mathrm{Re}_{\infty}^{-3/4}L_{\infty}$. Note that $\mathrm{Re}_{\infty}=u_{iso}L_{\infty}/\nu$ is the turbulent Reynolds number evaluated at the bottom of the surface-influenced region (at a distance of $L_{\infty}$ from the free surface) The data of CM03 suggested that the wall-parallel vorticity magnitude decays to zero within the viscous sublayer, whereas the wall-normal velocity magnitude grows linearly within the Kolmogorov sublayer. Note that the LES of CM03 with a Reynolds number of $\mathrm{Re}_{\tau}=1280$ was performed in a comparatively small computational domain ($L_{x}\leq 2\pi h$, $L_{z}\leq\pi h$) that is too restrictive to accommodate VLSMs [7]. Moreover, uncertainties from their sub-grid scale model motivates a follow-up investigation by means of DNS.
@@ 2 Numerical Method
Table 1: Turbulent channel flow simulation cases as published by [6, 7]. $L_{x}$ and $L_{z}$ are the stream- and spanwise domain length. $N_{x}$, $N_{y}$ and $N_{z}$ are the number of grid points with respect to the streamwise, wall-normal and radial direction; $\Delta x^{+}$, the streamwise grid spacing; $\Delta z^{+}$, the spanwise grid spacing; $\Delta y^{+}_{max}$, the maximum grid spacing in the wall-normal direction. Leading “O” indicates OCF cases, “C” CCF cases. $\Delta T$ is the temporal averaging interval for turbulent statistics.
case $\mathrm{Re}_{\tau}$ $\mathrm{Re}_{b}$ $\nicefrac{{L_{x}}}{{h}}$ $\nicefrac{{L_{z}}}{{h}}$ $N_{x}$ $N_{y}$ $N_{z}$ $\Delta x^{+}$ $\Delta z^{+}$ $\Delta y^{+}_{max}$ $\Delta Tu_{b}/h$
@@
Hereinafter, angular brackets denote averaging over both homogeneous directions and time. Normalization in viscous or “wall units” is indicated by a $+$ superscript. Corresponding scales are the friction velocity $u_{\tau}$ and the viscous length scale $\delta_{\nu}=\nu/u_{\tau}$. Characteristic scales for normalization in “bulk units” on the other hand are the bulk velocity $u_{b}$ and the channel height $h$. Hence, the bulk Reynolds number is defined as $\mathrm{Re}_{b}=u_{b}h/\nu$.
@@ 3 Results
As stated in section 1, CM03 defined the near-surface length scales based on a turbulent Reynolds number $\mathrm{Re}_{\infty}$ which characterizes the “turbulence below the surface-influenced region” (CM03). However, since in channel flow $\mathrm{Re}_{\infty}$ is closely linked to the bulk Reynolds number $\mathrm{Re}_{b}$, herein we define the near-surface length scales based on $\mathrm{Re}_{b}$ and on the channel height $h$, which are readily accessible parameters in wall-bounded turbulence. Henceforth we use the following definitions:
$\ell_{V}=\mathrm{Re}_{b}^{-1/2}h,$ (1)
$\ell_{K}=\mathrm{Re}_{b}^{-3/4}h,$ (2)
as characteristic length scales for the wall-normal extent of the near-surface viscous and Kolmogorov sublayers, respectively.
@@ 3.1 Mean velocity scaling
Figure 2: Surface-normal derivative of the mean streamwise velocity profile $d\langle u\rangle^{+}/d(y/h)$ in the vicinity of the free surface. Solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900. The distance from the free surface is normalized with (a) the channel height $h$; (b) the near-surface viscous length scale $\ell_{V}$; (c) the near-surface Kolmogorov length scale $\ell_{K}$.
Here, the distance from the free surface is normalized with different length scales, i.e. the channel height $h$ (figure 2a), the near-surface viscous length scale $\ell_{V}$ (figure 2b), and the near-surface Kolmogorov length scale $\ell_{K}$ (figure 2c). It can be seen that the streamwise velocity derivative profiles in the immediate vicinity of the free surface collapse well when the surface-normal coordinate is normalized with the near-surface Kolmogorov length scale $\ell_{K}$ (figure 2c). Moreover, the layer thickness where the mean velocity derivative rapidly decays to zero measures approximately $\delta_{K}\approx 20\ell_{K}$. Note that unlike anticipated by [18], a region where $\,{\rm d}\langle u\rangle/\,{\rm d}y$ stays approximately zero when moving away from free surface does not exist. Finally, the observed scaling clearly shows a necessity for fine grid resolutions near the free surface when performing OCF DNSs. We will return to this topic in section 3.5
@@ 3.2 Turbulence intensity scaling
As already discussed above, previous authors have defined a “surface-influenced layer” as the region where the mean-velocity profiles in OCF deviate from their CCF counterparts (CM03). Figure 3 confirms that the thickness of this layer scales with the channel height $h$, and that its extent can be quantified as $\delta_{NVD}\approx 0.3h$. Note that we deliberately use the new symbol $\delta_{NVD}$ (for “normal velocity damping layer”) instead of the denomination “surface-influenced layer” used by others, as the latter region will be shown to extend much further than previously assumed. As a consequence the commonly invoked three-layer structure of OCF will be extended to a four-layer structure.
@@
Figure 6: Turbulence intensities normalized by $u_{\tau}$ (a,c) and $u_{b}$ (b,d) as function of the distance from the wall $y/h$ for open channel (a,b) and closed channel configuration (c,d): , $u_{rms}$; , $v_{rms}$; , $w_{rms}$. (a,b) solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900 [7]. (c,d) solid lines, C200; dashed lines, C400; dashed-dotted lines, C600; dotted lines, C900 [7]. The insets in (b,d) show a zoom for the streamwise turbulence intensity component. The gray lines in (b) indicate OCF DNS data from [13] at $\mathrm{Re}_{\tau}=550$ ( ) and from Pirozzoli [38] at $\mathrm{Re}_{\tau}=1000$ (      ), 2000 (     ), 3000 (       ), 6000 (   ). The symbols ( ) in (b) indicate a profile from OCF measurements by Duan et al. [9, 12] at $\mathrm{Re}_{\tau}=2407$. The gray lines in (d) indicate CCF DNS data at $\mathrm{Re}_{\tau}=2003$ [ , 15], $\mathrm{Re}_{\tau}=5186$ [      , 24], $\mathrm{Re}_{\tau}=10049$ [     , 33].
@@
While the spanwise and wall-normal contributions to the turbulence intensity scale with the friction velocity $u_{\tau}$ in the bulk region of the flow for $\mathrm{Re}_{\tau}\geq 400$ for both flow configurations (figure 6a,c), the streamwise contribution differs. Figure 6(b) shows that the streamwise turbulence intensity in open channel flow appears to scale with the bulk velocity $u_{b}$ for $\mathrm{Re}_{\tau}\geq 400$ and $y/h\gtrsim 0.3$, whereas in the closed channel case it neither scales with $u_{\tau}$ nor with $u_{b}$ (figure 6c,d), as reported in Bauer et al. [7]. Note that the described scaling behavior is only observed for domain sizes $L_{x}\geq 12\pi h$ [7]. This explains why data at high Reynolds number computed in short computational domains [e.g. 38, cf. their figure 5] does not exhibit the scaling of $u_{rms}$ with $u_{b}$ observed here.
@@
Next we turn to the scaling of the turbulence intensities in the different surface layers in OCF. Figure 7 shows turbulence intensity profiles normalized in wall units, plotted as a function of the distance from the free surface normalized with various characteristic length scales ($h$, $\ell_{V}$, $\ell_{K}$).
@@
Figure 7: Turbulence intensities normalized in wall units as function of distance from free surface: , $u^{+}_{rms}$; , $v^{+}_{rms}$; , $w^{+}_{rms}$; solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900. The distance from the free surface is normalized with (a) the channel height $h$; (b) the near-surface viscous length scale $\ell_{V}$; (c) the near-surface Kolmogorov length scale $\ell_{K}$.
@@
It can be seen that scaling $y$ with $\ell_{V}$ provides the best collapse of the wall-normal component $v^{+}_{rms}$ which features a steep drop to zero within the thin viscous sublayer. The surface-normal derivative of the surface normal turbulence intensity is presented in figure 8. In the immediate vicinity of the free surface ($(h-y)/\ell_{V}\lesssim 1$) the wall-normal velocity grows linearly when moving away from the free surface, with a slope of approximately 0.09. Therefore, the viscous sublayer is estimated as $\delta_{V}\approx\ell_{V}$.
@@
Figure 8: Surface-normal derivative of the surface-normal turbulence intensity in the vicinity of the free surface $-dv^{+}_{rms}/d(y/\ell_{V})$. Solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900.
@@
Although the region where $\partial v_{rms}/\partial y$ is constant is often referred to as the Kolmogorov sublayer [4, CM03], we find that the layer scales clearly with the near-surface viscous scale $\ell_{V}$ rather than with the Kolmogorov scale $\ell_{K}$ (compare figure 7b and c).

### v2

@@ Abstract
Turbulent open channel flow is known to feature a multi-layer structure near the free surface. In the present work we employ direct numerical simulations considering Reynolds numbers up to $\mathrm{Re}_{\tau}=900$ and domain sizes large enough ($L_{x}=12\pi h$, $L_{z}=4\pi h$) to faithfully capture the effect of very-large-scale motions in order to test the proposed scaling laws and ultimately answer the question: How far does the influence of the free surface extend? In the region near the free surface, where fluctuation intensities of velocity and vorticity become highly anisotropic, we observe the previously documented triple-layer structure, consisting of a wall-normal velocity damping layer that scales with the channel height $h$, and two sublayers that scale with the near-surface viscous length scale $\ell_{\mathrm{V}}=\mathrm{Re}_{\mathrm{b}}^{-1/2}h$ and with the Kolmogorov length scale $\ell_{\mathrm{K}}=\mathrm{Re}_{\mathrm{b}}^{-3/4}h$, respectively. The scaling laws previously proposed by Calmet and Magnaudet [J. Fluid. Mech. 474, 355–378 (2003)] are found to hold with the following exceptions. The thin layer, where the intensity of surface-parallel components of the vorticity rapidly decreases to zero, is here found to scale with the Kolmogorov length scale $\ell_{\mathrm{K}}$ rather than with the near-surface viscous scale $\ell_{\mathrm{V}}$. In addition, we argue that the Kolmogorov length scale is the relevant scale for the mean velocity gradient near the free surface. Both the mean velocity gradient and the fluctuation intensity of the surface-parallel component of vorticity decay to zero in the Kolmogorov sublayer $\delta_{\mathrm{K}}\approx 20\ell_{\mathrm{K}}$. On the other hand, the layer, where the wall-normal turbulence intensity decreases linearly to zero near the free surface, scales with $\ell_{\mathrm{V}}$ rather than $\ell_{\mathrm{K}}$ as suggested by Calmet and Magnaudet. The corresponding near-surface viscous sublayer measures $\delta_{\mathrm{V}}\approx\ell_{\mathrm{V}}$. Importantly, the streamwise turbulence intensity profile for $\mathrm{Re}_{\tau}\geq 400$ suggests that the influence of the free-slip boundary penetrates essentially all the way down to the solid wall through the appearance of enhanced very-large-scale motions ($\delta_{\mathrm{SIL}}\approx h$). In contrast, the layer where the surface-normal turbulence intensity is damped to zero is restricted to the free surface ($\delta_{\mathrm{NVD}}\approx 0.3h$). As a consequence, the partitioning of the surface-influenced region has to be expanded to a four-layer structure that spans the entire channel height $h$.
@@ 1 Introduction
Plane Poiseuille flow, also known as closed channel flow (CCF), is one of the most studied canonical flows by means of direct numerical simulations [DNSs, 22, 27, 15, 24, 33]. The numerical domain is defined by doubly-periodic boundaries in the stream- and spanwise directions, and impermeable no-slip walls at the bottom and the top. Conversely, less attention has been paid to open channel flow (OCF), where one of the no-slip walls is replaced by a free-slip plane, despite its direct relevance to environmental flows. As noted by Pinelli et al. [37], OCF is of particular interest when studying the mass transfer across gas-liquid interfaces in oceans or engineering applications. One of the defining characteristics of OCF is the transfer of kinetic energy from the vertical to the surface-parallel velocity components due to the damping effect of the free surface [39, 25, 17, 23, 35]. While the inter-component kinetic energy transfer is dominated by the small scales in the vicinity of the free surface, Pinelli et al. [37] showed that low-speed large- and very-large-scale motions ((V)LSMs) are much more correlated with the regions of small-scale vortices than high-speed (V)LSMs. Hence, the spatial distribution of the small scales in the vicinity of the free surface is modulated by the large scales. Most of the early OCF DNSs were restricted to small friction Reynolds numbers $\mathrm{Re}_{\tau}=u_{\tau}h/\nu$, based on friction velocity $u_{\tau}$, channel height $h$ and kinematic viscosity $\nu$, and/or small computational domains. Recently, Yao et al. [41] compared CCF and OCF by means of DNS data in computational domains of $L_{x}\times L_{z}=8\pi h\times 4\pi h$ for Reynolds numbers up to $\mathrm{Re}_{\tau}=2000$. In agreement with experiments [9, 12] and previous DNSs [3], they reported that VLSMs—which are a common feature of wall-bounded turbulent flows—are more energetic and larger in OCF than in CCF. Generally, VLSMs become more energetic with increasing Reynolds number [3, 12, 41], leading to the failure of wall scaling of the streamwise intensity in CCF [15]. As reported by [15] for CCF, the streamwise turbulence intensity scales neither in wall nor in bulk units. To compare wall-attached motions between OCF and CCF, Gong et al. [13] performed DNSs in extremely large domains of length $48\pi h$ for $\mathrm{Re}_{\tau}=550$ and length $24\pi h$ for $\mathrm{Re}_{\tau}=1000$. They found that wall-attached motions contribute more to the turbulent intensity in OCF than in CCF. In Bauer et al. [7], the current authors carried out DNSs at friction Reynolds numbers up to $\mathrm{Re}_{\tau}\approx 900$ in large computational domains up to $L_{x}\times L_{z}=12\pi h\times 4\pi h$ and found that unlike CCF, the streamwise turbulence intensity in OCF scales with the bulk velocity for $\mathrm{Re}_{\tau}\gtrsim 400$. The additional streamwise kinetic energy in OCF was attributed to larger and more intense VLSMs compared to CCF.
@@
CM03 first deduced an integral length scale $L_{\infty}$ using the assumption of high-Reynolds-number equilibrium between kinetic energy transfer and dissipation rate, $\varepsilon\sim u_{iso}^{3}/\ell$, (where $u_{iso}=(2k/3)^{1/2}$, $k=\langle u_{i}^{\prime}u_{i}^{\prime}\rangle/2$ is the turbulent kinetic energy, and $\ell$ is the characteristic scale of the energetic eddies) and the relation $L_{\infty}=\ell/2$ which Hunt and Graham [14] found to be valid for homogeneous-isotropic turbulence. Subsequently, CM03 performed a large-eddy simulation (LES) of OCF at $\mathrm{Re}_{\tau}=1280$ and compared their results to these scaling relations. Based on their data, CM03 estimated the thickness of the so-called “surface-influenced layer” as $\delta_{\mathrm{SIL}}\approx L_{\infty}\approx 0.2h$, in agreement with the rapid distortion theory of Hunt and Graham [14]. CM03 further categorized the surface-influenced layer into the following sublayers in the near-surface region: a viscous sublayer (of wall-normal extent $\delta_{\mathrm{V}}$) which scales with $\mathrm{Re}_{\infty}^{-1/2}L_{\infty}$, and a Kolmogorov sublayer (of extent $\delta_{K}$) which scales with $\mathrm{Re}_{\infty}^{-3/4}L_{\infty}$. Note that $\mathrm{Re}_{\infty}=u_{iso}L_{\infty}/\nu$ is the turbulent Reynolds number evaluated at the bottom of the surface-influenced region (at a distance of $L_{\infty}$ from the free surface). The data of CM03 suggested that the wall-parallel vorticity magnitude decays to zero within the viscous sublayer, whereas the wall-normal velocity magnitude grows linearly within the Kolmogorov sublayer. Note that the LES of CM03 with a Reynolds number of $\mathrm{Re}_{\tau}=1280$ was performed in a comparatively small computational domain ($L_{x}\leq 2\pi h$, $L_{z}\leq\pi h$) that is too restrictive to accommodate VLSMs [7]. Moreover, uncertainties from their sub-grid scale model motivates a follow-up investigation by means of DNS.
@@ 2 Numerical Method
Table 1: Turbulent channel flow simulation cases as published by [6, 7]. $L_{x}$ and $L_{z}$ are the stream- and spanwise domain length. $N_{x}$, $N_{y}$ and $N_{z}$ are the number of grid points with respect to the streamwise, wall-normal and radial direction; $\Delta x^{+}$, the streamwise grid spacing; $\Delta z^{+}$, the spanwise grid spacing; $\Delta y^{+}_{\mathrm{max}}$, the maximum grid spacing in the wall-normal direction. Leading “O” indicates OCF cases, “C” CCF cases. $\Delta T$ is the temporal averaging interval for turbulent statistics.
case $\mathrm{Re}_{\tau}$ $\mathrm{Re}_{\mathrm{b}}$ $\nicefrac{{L_{x}}}{{h}}$ $\nicefrac{{L_{z}}}{{h}}$ $N_{x}$ $N_{y}$ $N_{z}$ $\Delta x^{+}$ $\Delta z^{+}$ $\Delta y^{+}_{\mathrm{max}}$ $\Delta Tu_{\mathrm{b}}/h$
@@
Hereinafter, angular brackets denote averaging over both homogeneous directions and time. Normalization in viscous or “wall units” is indicated by a $+$ superscript. Corresponding scales are the friction velocity $u_{\tau}$ and the viscous length scale $\delta_{\nu}=\nu/u_{\tau}$. Characteristic scales for normalization in “bulk units” on the other hand are the bulk velocity $u_{\mathrm{b}}$ and the channel height $h$. Hence, the bulk Reynolds number is defined as $\mathrm{Re}_{\mathrm{b}}=u_{\mathrm{b}}h/\nu$.
@@ 3 Results
As stated in section 1, CM03 defined the near-surface length scales based on a turbulent Reynolds number $\mathrm{Re}_{\infty}$ which characterizes the “turbulence below the surface-influenced region” (CM03). However, since in channel flow $\mathrm{Re}_{\infty}$ is closely linked to the bulk Reynolds number $\mathrm{Re}_{\mathrm{b}}$, herein we define the near-surface length scales based on $\mathrm{Re}_{\mathrm{b}}$ and on the channel height $h$, which are readily accessible parameters in wall-bounded turbulence. $\mathrm{Re}_{\infty}$, on the other hand, is computed from $u_{iso}$ at the bottom of the surface-influenced layer, which supposedly corresponds to the most isotropic region of the flow. However, our OCF data show that the flow becomes less isotropic at higher Reynolds numbers due to VLSMs, which weakens the isotropy assumption. Moreover, the choice of the wall-normal location where $u_{iso}$ is computed is not trivial, as for instance the wall-normal coordinate of the most isotropic region might vary with the Reynolds number. Consequently, in OCF $\mathrm{Re}_{\mathrm{b}}$ is a more robust measure than $\mathrm{Re}_{\infty}$, since it is an integral quantity. Henceforth we use the following definitions:
$\ell_{\mathrm{V}}=\mathrm{Re}_{\mathrm{b}}^{-1/2}h,$ (1)
$\ell_{\mathrm{K}}=\mathrm{Re}_{\mathrm{b}}^{-3/4}h,$ (2)
as characteristic length scales for the wall-normal extent of the near-surface viscous and Kolmogorov sublayers, respectively. Note that what is called the viscous sublayer near the free surface differs from Prandtl’s viscous sublayer near the solid wall. Although viscous effects are much weaker near the free surface than near the solid wall, they also play a significant role in the vicinity of the free surface.
@@ 3.1 Mean velocity scaling
Figure 2: Surface-normal derivative of the mean streamwise velocity profile $d\langle u\rangle^{+}/d(y/h)$ in the vicinity of the free surface. Solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900. The distance from the free surface is normalized with (a) the channel height $h$; (b) the near-surface viscous length scale $\ell_{\mathrm{V}}$; (c) the near-surface Kolmogorov length scale $\ell_{\mathrm{K}}$.
Here, the distance from the free surface is normalized with different length scales, i.e. the channel height $h$ (figure 2a), the near-surface viscous length scale $\ell_{\mathrm{V}}$ (figure 2b), and the near-surface Kolmogorov length scale $\ell_{\mathrm{K}}$ (figure 2c). It can be seen that the streamwise velocity derivative profiles in the immediate vicinity of the free surface collapse well when the surface-normal coordinate is normalized with the near-surface Kolmogorov length scale $\ell_{\mathrm{K}}$ (figure 2c). Moreover, the layer thickness where the mean velocity derivative rapidly decays to zero measures approximately $\delta_{\mathrm{K}}\approx 20\ell_{\mathrm{K}}$. Note that unlike anticipated by [18], a region where $\,{\rm d}\langle u\rangle/\,{\rm d}y$ stays approximately zero when moving away from free surface does not exist. Finally, the observed scaling clearly shows a necessity for fine grid resolutions near the free surface when performing OCF DNSs. We will return to this topic in section 3.5
@@ 3.2 Turbulence intensity scaling
As already discussed above, previous authors have defined a “surface-influenced layer” as the region where the mean-velocity profiles in OCF deviate from their CCF counterparts (CM03). Figure 3 confirms that the thickness of this layer scales with the channel height $h$, and that its extent can be quantified as $\delta_{\mathrm{NVD}}\approx 0.3h$. Note that we deliberately use the new symbol $\delta_{\mathrm{NVD}}$ (for “normal velocity damping layer”) instead of the denomination “surface-influenced layer” used by others, as the latter region will be shown to extend much further than previously assumed. As a consequence the commonly invoked three-layer structure of OCF will be extended to a four-layer structure.
@@
Figure 6: Turbulence intensities normalized by $u_{\tau}$ (a,c) and $u_{\mathrm{b}}$ (b,d) as function of the distance from the wall $y/h$ for open channel (a,b) and closed channel configuration (c,d): , $u_{rms}$; , $v_{rms}$; , $w_{rms}$. (a,b) solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900 [7]. (c,d) solid lines, C200; dashed lines, C400; dashed-dotted lines, C600; dotted lines, C900 [7]. The insets in (b,d) show a zoom for the streamwise turbulence intensity component. The gray lines in (b) indicate OCF DNS data from [13] at $\mathrm{Re}_{\tau}=550$ ( ) and from Pirozzoli [38] at $\mathrm{Re}_{\tau}=1000$ (      ), 2000 (     ), 3000 (       ), 6000 (   ). The symbols ( ) in (b) indicate a profile from OCF measurements by Duan et al. [9, 12] at $\mathrm{Re}_{\tau}=2407$. The gray lines in (d) indicate CCF DNS data at $\mathrm{Re}_{\tau}=2003$ [ , 15], $\mathrm{Re}_{\tau}=5186$ [      , 24], $\mathrm{Re}_{\tau}=10049$ [     , 33].
@@
While the spanwise and wall-normal contributions to the turbulence intensity scale with the friction velocity $u_{\tau}$ in the bulk region of the flow for $\mathrm{Re}_{\tau}\geq 400$ for both flow configurations (figure 6a,c), the streamwise contribution differs. Figure 6(b) shows that the streamwise turbulence intensity in open channel flow appears to scale with the bulk velocity $u_{\mathrm{b}}$ for $\mathrm{Re}_{\tau}\geq 400$ and $y/h\gtrsim 0.3$, whereas in the closed channel case it neither scales with $u_{\tau}$ nor with $u_{\mathrm{b}}$ (figure 6c,d), as reported in Bauer et al. [7]. Note that the described scaling behavior is only observed for domain sizes $L_{x}\geq 12\pi h$ [7]. This explains why data at high Reynolds number computed in short computational domains [e.g. 38, cf. their figure 5] does not exhibit the scaling of $u_{rms}$ with $u_{\mathrm{b}}$ observed here.
@@
Next we turn to the scaling of the turbulence intensities in the different surface layers in OCF. Figure 7 shows turbulence intensity profiles normalized in wall units, plotted as a function of the distance from the free surface normalized with various characteristic length scales ($h$, $\ell_{\mathrm{V}}$, $\ell_{\mathrm{K}}$).
@@
Figure 7: Turbulence intensities normalized in wall units as function of distance from free surface: , $u^{+}_{rms}$; , $v^{+}_{rms}$; , $w^{+}_{rms}$; solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900. The distance from the free surface is normalized with (a) the channel height $h$; (b) the near-surface viscous length scale $\ell_{\mathrm{V}}$; (c) the near-surface Kolmogorov length scale $\ell_{\mathrm{K}}$.
@@
It can be seen that scaling $y$ with $\ell_{\mathrm{V}}$ provides the best collapse of the wall-normal component $v^{+}_{rms}$ which features a steep drop to zero within the thin viscous sublayer. The surface-normal derivative of the surface normal turbulence intensity is presented in figure 8. In the immediate vicinity of the free surface ($(h-y)/\ell_{\mathrm{V}}\lesssim 1$) the wall-normal velocity grows linearly when moving away from the free surface, with a slope of approximately 0.09. Therefore, the viscous sublayer is estimated as $\delta_{\mathrm{V}}\approx\ell_{\mathrm{V}}$.
@@
Figure 8: Surface-normal derivative of the surface-normal turbulence intensity in the vicinity of the free surface $-dv^{+}_{rms}/d(y/\ell_{\mathrm{V}})$. Solid lines, O200; dashed lines, O400; dashed-dotted lines, O600; dotted lines, O900.
@@
Although the region where $\partial v_{rms}/\partial y$ is constant is often referred to as the Kolmogorov sublayer [4, CM03], we find that the layer scales clearly with the near-surface viscous scale $\ell_{\mathrm{V}}$ rather than with the Kolmogorov scale $\ell_{\mathrm{K}}$ (compare figure 7b and c).